OpenStudy (anonymous):

Suppose f(x) = 2x - 1 and g(x) = x^2. Find (f+g)(x) and state domain.

jimthompson5910 (jim_thompson5910):

(f+g)(x) is the same as f(x) + g(x) so you just add up the functions and combine like terms if possible

OpenStudy (anonymous):

x^2 + 2x - 1 I get that part, it's more so that I have trouble understanding how to find domain.

jimthompson5910 (jim_thompson5910):

Rule: the domain of ANY polynomial is the set of all real numbers. You can plug in any real number in for x, for any polynomial, and you'll get some real number out.

jimthompson5910 (jim_thompson5910):

There are no restrictions (eg: division by zero errors) to worry about

OpenStudy (anonymous):

Okay! Thank you. Can you help me w/ another problem?

jimthompson5910 (jim_thompson5910):

Sure, go ahead

OpenStudy (anonymous):

Find h(k(x)) and k(h(x)), and state the domain. h(x) = x^2 - 3x + 1 and k(x) = log5(x). So, \[(\log_{5} )^{2} - 3(\log_{5} ) + 1\]

jimthompson5910 (jim_thompson5910):

I think you meant to say \[\Large h(k(x)) = \left(\log_{5}(x)\right)^2 + 3\log_{5}(x) + 1\] right?

OpenStudy (anonymous):

Yeah, I forgot the x's.

jimthompson5910 (jim_thompson5910):

as for the domain of h(k(x)), any ideas?

OpenStudy (anonymous):

\[[1, \infty)\]

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

you can plug in numbers for x smaller than 1 like x = 0.5 but you cannot plug in x = 0 or anything negative

jimthompson5910 (jim_thompson5910):

so the domain is x > 0 which in interval notation is \(\Large (0,\infty)\)

OpenStudy (anonymous):

Okay, that makes sense. Thanks!

jimthompson5910 (jim_thompson5910):

how about k(h(x)) ?

OpenStudy (anonymous):

Would log5(x^2 - 3x + 1)'s domain be (-inf, +inf)?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

hint: look at the graph of y = x^2 - 3x + 1. Are there points on that parabola that have y coordinates that are either 0 or a negative number?

OpenStudy (anonymous):

Yes, (1, -1) and (2, -1).

OpenStudy (anonymous):

Well, looking at the table tells me that.

jimthompson5910 (jim_thompson5910):

so because of that, log(x^2 - 3x + 1) = log(-1) when x = 1 but log(-1) leads to a non-real number. So x = 1 is NOT part of the domain

jimthompson5910 (jim_thompson5910):

there are other x values that aren't part of the domain for this reason

OpenStudy (anonymous):

Would it be\[(-\infty, 1] \cup [1, 2] \cup [2, \infty)\]

jimthompson5910 (jim_thompson5910):

what are the roots of x^2 - 3x + 1 ?

OpenStudy (anonymous):

Are there even any? Doing the big x, what is there that multiplies to 1 and adds to three?

jimthompson5910 (jim_thompson5910):

no, so you'll need to use the quadratic formula

jimthompson5910 (jim_thompson5910):

x^2 - 3x + 1 does not factor the roots are decimal values

OpenStudy (anonymous):

5/2 and 1/2?

jimthompson5910 (jim_thompson5910):

I graphed x^2 - 3x + 1 with a graphing calculator https://www.desmos.com/calculator/fp80fcskll click on the x-intercepts and the coordinates will pop up. Tell me what values pop up

OpenStudy (anonymous):

0.382 2.618 I'm so mad b/c I got those before but thought they were wrong.

jimthompson5910 (jim_thompson5910):

notice how the portion between 0.382 and 2.618 is below the x axis the points on the parabola in this interval have negative y coordinates or the y coordinates are 0. So we have to kick this interval out of the domain Start with \(\Large (-\infty, \infty)\) and kick out that interval to end up with \(\Large (-\infty, 0.382)\cup(2.618, \infty)\)

jimthompson5910 (jim_thompson5910):

So the domain of log(x^2 - 3x + 1) is \(\Large (-\infty, 0.382)\cup(2.618, \infty)\)

OpenStudy (anonymous):

Thank you for all of your help. I think I understand what to do now.

jimthompson5910 (jim_thompson5910):

sure thing, glad to be of help